# Length independent generalization bounds for deep SSM architectures with stability constraints

### v1

@@ Appendix B Proofs
The restriction of an MLP to the ball $B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r)$ maps to the ball $B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{v}})}(K_{W}r+K_{\mathbf{b}})$, because the elementwise ReLU does not increase the infinity norm, hence we can apply Lemma 21.
@@
Proof [Proof of Theorem 18] We wish to apply Theorem 20 to the set of deep SSM models $\mathcal{F}$. Let us fix a random sample $S=\{\mathbf{u}_{1},\ldots,\mathbf{u}_{N}\}\subset\left(\ell^{2,2}(\mathbb{R}^{n_{\text{in}}})\right)^{N}$. As the loss function is Lipschitz according to Assumption 9, we have that for any $f\in\mathcal{F}$
$|l(f(\mathbf{u}),y)|\leq 2L_{l}\max\{f(\mathbf{u}),y\}\leq 2L_{l}\max\{K_{\text{Dec}}r_{L+2},K_{y}\},$
thus $K_{l}\leq 2L_{l}\max\{K_{\text{Dec}}r_{L+2},K_{y}\}$.
@@
Therefore, the SSM layer in the first SSM block is considered as a map $B_{\ell^{2,2}(\mathbb{R}^{n_{u}})}(K_{\text{Enc}}K_{\mathbf{u}})\to B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r_{1})$, while the rest of the SSM layers in the SSM blocks are considered as a map $B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r_{i})\to B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r_{i+1})$.This is needed, because the Encoder is constant in time, therefore the Composition Lemma wouldn’t be able to carry the $\ell^{2,2}$ norm of the input through the chain of estimation along the entire model. This is one of the key technical points which makes it possible to establish a time independent bound.
By the conditions of the Theorem and the stability assumption in Assumption 9 we have that the Encoder, Decoder, Pooling, SSM and MLP layers are each $(\mu,c)$-RC for some $\mu$ and $c$ from Lemma 13. By Lemma 11 we have that the composition of an SSM layer and an MLP is $(\mu,c)$-RC. A residual SSM block is then $(\mu+\alpha,c)$-RC, because
$\mathbb{E}_{\sigma}\left[\sup\limits_{g\circ\mathcal{S}_{\Sigma}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}(g(\mathcal{S}_{\Sigma}(\mathbf{z}_{i}))+\alpha\mathbf{z}_{i})\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]\leq$
$\mathbb{E}_{\sigma}\left[\sup\limits_{g\circ\mathcal{S}_{\Sigma}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}g(\mathcal{S}_{\Sigma}(\mathbf{z}_{i}))\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]+\alpha\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
$\leq(\mu+\alpha)\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]+\frac{c}{\sqrt{N}}$
Hence, by Corollary 12, the whole deep SSM model is $(\mu,c)$-RC as a map between $X_{1}=B_{\ell^{2,2}(\mathbb{R}^{n_{\text{in}}})}(K_{\mathbf{u}})$ and $X_{2}=(\mathbb{R},|\cdot|)$. The Theorem is then a direct corollary of the following Lemma.
Lemma 24
Let $\mathcal{F}$ be a set of functions that is $(\mu,c)$-RC w.r.t. $X_{1}=B_{\ell^{2,2}(\mathbb{R}^{n_{\text{in}}})}(K_{\mathbf{u}})$ and $X_{2}=(\mathbb{R},|\cdot|)$. The Rademacher complexity of $\mathcal{F}$ w.r.t. some dataset $S$ for which Assumption 9 holds, admits the following inequality.
@@
$\leq\mu\mathbb{E}_{\sigma}\left[\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right\rVert_{\ell^{2,2}(\mathbb{R}^{n_{\text{in}}})}\right]+\frac{c}{\sqrt{N}}$
@@
$\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right\rVert_{\ell^{2,2}(\mathbb{R}^{n_{\text{in}}})}=\sqrt{\sum\limits_{k=1}^{T}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}[k]\right\rVert^{2}_{2}}$
@@
$\mathbb{E}_{\sigma}\left[\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right\rVert_{\ell^{2,2}(\mathbb{R}^{n_{\text{in}}})}\right]=\mathbb{E}_{\sigma}\left[\sqrt{\sum\limits_{k=1}^{T}\frac{1}{N^{2}}\sum\limits_{i=1}^{N}\sum\limits_{j=1}^{N}\sigma_{i}\sigma_{j}\left\langle\mathbf{u}_{i}[k],\mathbf{u}_{j}[k]\right\rangle_{\mathbb{R}^{n_{\text{in}}}}}\right]$
@@
$=\sqrt{\frac{1}{N^{2}}\sum\limits_{i=1}^{N}\sum\limits_{k=1}^{T}\left\lVert\mathbf{u}_{i}[k]\right\rVert_{2}^{2}}=\sqrt{\frac{1}{N^{2}}\sum\limits_{i=1}^{N}\left\lVert\mathbf{u}_{i}\right\rVert^{2}_{\ell^{2,2}(\mathbb{R}^{n_{\text{in}}})}}\leq\sqrt{\frac{1}{N^{2}}NK_{\mathbf{u}}^{2}}\leq\frac{K_{\mathbf{u}}}{\sqrt{N}}$
@@
The constants $\mu$ and $c$ are obtained by substituting the results of Lemma 13 into the Corollary 12.

### v2

@@ D.1 Proofs for elements of deep SSMs
The restriction of an MLP to the ball $B_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r)$ maps to the ball $B_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{v}})}(K_{W}r+K_{\mathbf{b}})$, because the elementwise ReLU does not increase the infinity norm, hence we can apply Lemma D.1 and Remark D.2. Again, for the deep model the result is straightforward from Lemma 3.2 along with Lemma D.1, Remark D.2 and Remark E.1.
GLU. First of all, we show that the function $h:(\mathbb{R}^{2},\left\lVert\cdot\right\rVert_{2})\to(\mathbb{R},|\cdot|)$ defined as $h(\mathbf{x})=x_{1}\cdot\sigma(x_{2})$ is $\sqrt{2}(K+1)$-Lipschitz on a bounded domain, where $|x_{i}|\leq K$ for all $\mathbf{x}\in\mathbb{R}^{2}$ we consider. We will later specify the value of $K$ in relation to Assumption C.1. By the sigmoid being 1-Lipschitz, we have
$\leq\sqrt{2}(K+1)\left\lVert\mathbf{x}-\mathbf{y}\right\rVert_{2}$
Second, we recall Corollary 4 in Maurer (2016).
Theorem D.7 (Maurer (2016)).
Let $\mathcal{X}$ be any set, $(\mathbf{x}_{1},\ldots,\mathbf{x}_{N})\in\mathcal{X}^{N}$, let $\mathcal{F}$ be a set of functions $f:\mathcal{X}\to\ell_{T}^{2}(\mathbb{R}^{m})$ and let $h:\ell_{T}^{2}(\mathbb{R}^{m})\to\mathbb{R}$ be an $L$-Lipschitz function. Under $f_{k}$ denoting the $k$-th component function of $f$ and $\sigma_{ik}$ being a doubly indexed Rademacher variable, we have
We wish to apply Theorem D.7 to GLU layers. For any $Z\subseteq\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})$, by letting $GLU_{W}(\mathbf{z})=f_{GLU}(\mathbf{z})$ we have
$\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}GLU_{W}(\mathbf{z}_{i})\right\rVert_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
Now this is an alternative version of the Rademacher complexity, where we take the absolute value of the Rademacher average. In order to apply Theorem D.7, we reduce the problem to the usual Rademacher complexity. In turn, we can apply the last chain of inequalities in the proof of Proposition 6.2 in Hajek and Raginsky (2019). Concretely, by denoting $\mathbf{O}=\{\mathbf{0}\}_{i=1}^{N}$ and noticing that $GLU_{W}(0)=0$, we have
Let $\mathbf{x}_{i}=i$, $i=1,\ldots,N$ and let =$\mathcal{F}=\{f_{W,\underline{z},k,j}\mid(W,\underline{z},k,j)\in\mathcal{W}\times(Z\cup\{0\})\times[T]\times[n_{u}]\}$ such that $f_{W,\underline{z},k,j}(\mathbf{x}_{i})=\begin{bmatrix}GELU(\mathbf{z}_{i}[k])^{(j)}&(W(GELU(\mathbf{z}_{i}[k])))^{(j)}\end{bmatrix}^{T}$ for $\underline{z}=\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z$. Since $Z\subseteq(B_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}(K_{\mathbf{u}}))^{N}$, it follows for all $\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z$ and for all $k\in\mathbb{N}$ that $\|\mathbf{z}_{i}[k]\|_{\infty}\leq K_{\mathbf{u}}$, and hence $|GELU(\mathbf{z}_{i}[k])^{(j)}|<K_{\mathbf{u}}$, leading to $|W(GELU(\mathbf{z}_{i}[k]))^{(j)}|<\sup\limits_{W\in\mathcal{W}}\|W\|_{\infty,\infty}\cdot K_{\mathbf{u}}$. In particular, $GLU^{(j)}_{W}(\mathbf{z}_{i})[k]=h(f_{W,\underline{z},k,j}(\mathbf{x}_{i}))=h|_{B}(f_{W,\underline{z},k,j}(\mathbf{x}_{i}))$, where $h|_{B}$ is the restriction of $h$ to $B=\{x\in\mathbb{R}^{2}\mid\|x\|_{\infty}<K\}$, and hence $h|_{B}$ is $\sqrt{2}(K+1)$-Lipschitz. Therefore we can set $K=\max\{K_{\mathbf{u}},\sup\limits_{W\in\mathcal{W}}\|W\|_{\infty,\infty}\cdot K_{\mathbf{u}}\}$.
We are ready to apply Theorem D.7, together with the GLU definition and its $\sqrt{2}(K+1)$-Lipschitzness, we have
$=2\mathbb{E}_{\sigma}\left[\sup\limits_{f\in\mathcal{F}}\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}h(f(\mathbf{x}_{i}))\right]\leq 4(K+1)\underbrace{\mathbb{E}_{\sigma}\left[\sup\limits_{f\in\mathcal{F}}\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}GELU(\mathbf{z}_{i}[k])^{(j)}\right]}_{A}$
$A=\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z\cup\{\mathbf{O}\}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}GELU(\mathbf{z}_{i})\right\rVert_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]=$
$\leq 4\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z\cup\{\mathbf{O}\}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]=4\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
and
$B=\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in\{\mathbf{O}\}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}W(GELU(\mathbf{z}_{i}))\right\rVert_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
$\leq\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{\infty,\infty}\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z\{\mathbf{O}\}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}GELU(\mathbf{z}_{i})\right\rVert_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
$\leq 4\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{\infty,\infty}\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
Here we used the linearity of $W$ and the exact same calculation as in the proof of Lemma D.1.
By combining the inequalities above, it follows that
$16(K+1)\left(\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{\infty,\infty}+1\right)\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
Substituting the value of $K$ gives the result.
D.2 Proof of compostion Lemma 3.2
Proof D.8 (Proof of Lemma 3.2).
Let $Z\subseteq X_{1}^{N}$ and $\tilde{Z}=\{\{\varphi_{1}(\mathbf{u}_{i})\}_{i=1}^{N}\mid\varphi_{1}\in\Phi_{1}\}$. We have
The upcoming corollary is straightforward by induction along with the fact that the pooling layer $f^{\text{Pool}}$ is $(1,0)$-RC (see Lemma D.4).
Corollary D.9.
Let $\mathcal{F}$ be a set of deep SSM models, i.e. according to Definition of deep SSM, for $f\in\mathcal{F}$ we have $f=f^{\text{Dec}}\circ f^{\text{Pool}}\circ f^{\text{B}_{L}}\circ\ldots\circ f^{\text{B}_{1}}\circ f^{\text{Enc}}$, such that $f^{\text{Enc}}$, $f^{\text{Dec}}$ and each $f^{\text{B}_{i}}$ are from $(\mu_{0},c_{0})$-RC, $(\mu_{L+1},c_{L+1})$-RC and $(\mu_{i},c_{i})$-RC sets respectively for all $i$. Then $\mathcal{F}$ is $\left(\prod\limits_{i=0}^{L+1}\mu_{i},\sum\limits_{j=1}^{L}\left[\prod\limits_{i=j+1}^{L+1}\mu_{i}\right]c_{j}\right)$-RC.
We can see that the SSM layer can only increase the input’s complexity by the factor $\left\lVert\Sigma\right\rVert_{p}$, $p=1,2$, a quantity that gets smaller as the system gets more stable. This gets even more crucial when dealing with long range sequences, because the Neural Network layers are constant in time.
Remark D.10.
The results of of Lemma D.4 hold for unbounded input spaces as well. The reason for restricting the input space to a ball of radius $r$ is the composition with MLPs or GLU layers, as discussed in the interpretation of Definition 3.1.
Appendix E Statement and proof of the main theorem
So far we showed in D that each component of a deep SSM model satisfies Definition 3.1. We also proved that the composition of such components also satisfies the definition. The main theorem summarizes these results and exploits the fact that the Rademacher complexity of a $(\mu,c)$-RC set of models is upper bounded by terms containing $\mu$ and $c$.
Before we state the formal theorem let us discuss balls in Banach spaces regarding the contraction lemma.
Remark E.1.
The results of of Lemma D.4 hold for unbounded input spaces as well. The reason for restricting the input space to a ball of radius $r$ is the composition with GLU layers, as discussed in the interpretation of Definition 3.1. As a result of Lemma D.4, the restriction of a deep SSM model to a ball of an arbitrary radius $r$ has its image contain in a ball with a radius $\hat{r}$ depending on $r$ and the possible parameter set of each layer in the composite model. The exact value of $\hat{r}$ can be calculated using Lemma D.4. Namely, consider a residual SSM block, defined as $f^{\text{B}}(\mathbf{z})[k]=g(\mathcal{S}_{\Sigma_{i}}(\mathbf{z})[k])+\alpha\mathbf{z}[k]$ for all $k\in[T]$. Let $R_{g}(r)$ and $R_{\Sigma}(r)$ be the radius of the ball that is an image of the ball of radius $r$ in the domain, which can be obtained from Lemma D.4. Then the radius $\hat{r}=R_{f}(r)=(R_{g}(r)R_{\Sigma}(r)+\alpha)r$. We can apply this formula recursively to get the radius belonging to a deep SSM model containing several blocks.
Theorem E.2.
Let $\mathcal{F}$ be a set of deep SSM models, namely let $f\in\mathcal{F}$ has the form $f=f^{\text{Dec}}\circ f^{\text{Pool}}\circ f^{\text{B}_{L}}\circ\ldots\circ f^{\text{B}_{1}}\circ f^{\text{Enc}}$ with layer parameter sets $\mathcal{W}_{\text{Dec}}$, $\mathcal{W}_{\text{B}_{L}},\ldots,\mathcal{W}_{\text{B}_{1}}$ and $\mathcal{W}_{\text{Enc}}$ respectively, where $f^{\text{B}_{i}}$ is an SSM block for all $i$, i.e. $f^{\text{B}_{i}}(\mathbf{z})[k]=g_{i}(\mathcal{S}_{\Sigma_{i}}(\mathbf{z})[k])+\alpha_{i}\mathbf{z}[k]$ for all $k\in[T]$, and $\mathcal{W}_{\text{B}_{i}}=\mathcal{E}_{i}\times\mathcal{W}_{g_{i}}$. Let Assumption C.1 hold and let us assume that each $f\in\mathcal{F}$ maps from $B_{(\mathbb{R}^{n_{u}},\left\lVert\cdot\right\rVert_{2})}(K_{\mathbf{u}})$, and each set of nonlinearities $g_{i}$ is $(\mu_{g_{i}},c_{g_{i}})$-RC.
Under these assumptions, there exists $\hat{r}<\infty$ such that the image of each $f\in\mathcal{F}$ is contained in the ball $B_{(\mathbb{R},|\cdot|)}(\hat{r})$ and the following holds with probability at least $1-\delta$.
$\mu\leq K_{\text{Enc}}K_{\text{Dec}}\left(\mu_{g_{1}}K_{2}+\alpha_{1}\right)\prod\limits_{i=2}^{L}\left(\mu_{g_{i}}K_{1}+\alpha_{i}\right)$, $c\leq K_{\text{Dec}}\sum\limits_{j=1}^{L}\left[\prod\limits_{i=j+1}^{L}\left(\mu_{g_{i}}K_{1}+\alpha_{i}\right)\right]c_{g_{j}}$ and $K_{l}>0$ such that $|l(\cdot,\cdot)|\leq K_{l}$. In particular, we obtain $K_{l}\leq 2L_{l}\max\{K_{\text{Dec}}\hat{r},K_{y}\}$.
Proof E.3 (Proof of Theorem E.2).
First, let us recite the definition of Rademacher complexity.
Definition E.4.
The following is a standard theorem we use in the proof.
Theorem E.5.
@@ Appendix E Statement and proof of the main theorem
We wish to apply the Theorem E.5 to the set of deep SSM models $\mathcal{F}$. Let us fix a random sample $S=\{\mathbf{u}_{1},\ldots,\mathbf{u}_{N}\}\subset\left(\ell_{T}^{2,2}(\mathbb{R}^{n_{\text{in}}})\right)^{N}$. As the loss function is Lipschitz according to Assumption C.1, we have that for any $f\in\mathcal{F}$
$|l(f(\mathbf{u}),y)|\leq 2L_{l}\max\{f(\mathbf{u}),y\}\leq 2L_{l}\max\{K_{\text{Dec}}\hat{r},K_{y}\},$
thus $K_{l}\leq 2L_{l}\max\{K_{\text{Dec}}\hat{r}\,K_{y}\}$. The constant $\hat{r}$ exists as a corollary of Lemma D.4, see Remark E.1.
@@
where the constants $r_{i}$ exist as a corollary of Lemma D.4, see Remark E.1. Therefore, the SSM layer in the first SSM block is considered as a map $B_{\ell_{T}^{2,2}(\mathbb{R}^{n_{u}})}(K_{\text{Enc}}K_{\mathbf{u}})\to B_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r_{1})$, while the rest of the SSM layers in the SSM blocks are considered as a map $B_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r_{i})\to B_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r_{i+1})$. This is needed, because the Encoder is constant in time, therefore the Composition Lemma wouldn’t be able to carry the $\ell_{T}^{2,2}$ norm of the input through the chain of estimation along the entire model. This is one of the key technical points which makes it possible to establish a time independent bound.
By the conditions of the Theorem and the stability assumption in Assumption C.1 we have that the Encoder, Decoder, Pooling, SSM and MLP layers are each $(\mu,c)$-RC for some $\mu$ and $c$ from Lemma D.4. By Lemma 3.2 we have that the composition of an SSM layer and an MLP is $(\mu,c)$-RC. A residual SSM block is then $(\mu+\alpha,c)$-RC, because
$\mathbb{E}_{\sigma}\left[\sup\limits_{g\circ\mathcal{S}_{\Sigma}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}(g(\mathcal{S}_{\Sigma}(\mathbf{z}_{i}))+\alpha\mathbf{z}_{i})\right\rVert_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]\leq$
$\mathbb{E}_{\sigma}\left[\sup\limits_{g\circ\mathcal{S}_{\Sigma}}\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}g(\mathcal{S}_{\Sigma}(\mathbf{z}_{i}))\right\rVert_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]+\alpha\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]$
$\leq(\mu+\alpha)\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{z}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{z}_{i}\right\rVert_{\ell_{T}^{\infty,\infty}(\mathbb{R}^{n_{u}})}\right]+\frac{c}{\sqrt{N}}$
Hence, by Corollary D.9, the whole deep SSM model is $(\mu,c)$-RC as a map between $X_{1}=B_{\ell_{T}^{2,2}(\mathbb{R}^{n_{\text{in}}})}(K_{\mathbf{u}})$ and $X_{2}=(\mathbb{R},|\cdot|)$. The Theorem is then a direct corollary of the following Lemma.
Lemma E.6.
Let $\mathcal{F}$ be a set of functions between $X_{1}=B_{\ell_{T}^{2,2}(\mathbb{R}^{n_{\text{in}}})}(K_{\mathbf{u}})$ and $X_{2}=(\mathbb{R},|\cdot|)$ that is $(\mu,c)$-RC. The Rademacher complexity of $\mathcal{F}$ w.r.t. some dataset $S$ for which Assumption C.1 holds, admits the following inequality.
@@
$\leq\mathbb{E}_{\sigma}\left[\sup\limits_{f\in\mathcal{F}}\left|\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}f(\mathbf{u}_{i})\right|\right]\leq\mu\mathbb{E}_{\sigma}\left[\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right\rVert_{\ell_{T}^{2,2}(\mathbb{R}^{n_{\text{in}}})}\right]+\frac{c}{\sqrt{N}}$
@@
$\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right\rVert_{\ell_{T}^{2,2}(\mathbb{R}^{n_{\text{in}}})}=\sqrt{\sum\limits_{k=1}^{T}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}[k]\right\rVert^{2}_{2}}$
@@
$\mathbb{E}_{\sigma}\left[\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right\rVert_{\ell_{T}^{2,2}(\mathbb{R}^{n_{\text{in}}})}\right]=\mathbb{E}_{\sigma}\left[\sqrt{\sum\limits_{k=1}^{T}\frac{1}{N^{2}}\sum\limits_{i=1}^{N}\sum\limits_{j=1}^{N}\sigma_{i}\sigma_{j}\left\langle\mathbf{u}_{i}[k],\mathbf{u}_{j}[k]\right\rangle_{\mathbb{R}^{n_{\text{in}}}}}\right]$
@@
$=\sqrt{\frac{1}{N^{2}}\sum\limits_{i=1}^{N}\sum\limits_{k=1}^{T}\left\lVert\mathbf{u}_{i}[k]\right\rVert_{2}^{2}}=\sqrt{\frac{1}{N^{2}}\sum\limits_{i=1}^{N}\left\lVert\mathbf{u}_{i}\right\rVert^{2}_{\ell_{T}^{2,2}(\mathbb{R}^{n_{\text{in}}})}}\leq\sqrt{\frac{1}{N^{2}}NK_{\mathbf{u}}^{2}}\leq\frac{K_{\mathbf{u}}}{\sqrt{N}}$
@@
The constants $\mu$ and $c$ are obtained by substituting the results of Lemma D.4 into the Corollary D.9.